OpenStudy (anonymous):

write the first expression in terms of the second sec^2t x sin^2t, cos t

OpenStudy (anonymous):

This is the question stated in the correct format correct? \[1st = \sec ^{2t}(x) \sin ^{2t}(x)\] \[2nd \cos (t)\]

OpenStudy (anonymous):

im sorryi was trying to write\[\sec ^{2}t \sin ^{2}t , \cos t\]

OpenStudy (anonymous):

change sin^2(t) into [1- cos^2(t)] using --> identity: sin^2(t)+cos^2(t) = 1

OpenStudy (anonymous):

and the sec^2 (t) ? to 1/ cos^2(t)?

OpenStudy (anonymous):

yes, and then distribute

OpenStudy (anonymous):

so it would be [1- cos^2(t)] / cos^2 (t) ?

OpenStudy (anonymous):

im sorry im confused what am i distributing?

OpenStudy (anonymous):

hold on i need to do this on paper

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

[1- cos^2(t)] would change to sin^2(t) then that would be sin^2 (t) / cos^2 (t) = tan^2 (t)

OpenStudy (anonymous):

could you please rewrite the quesiton stating in full what it is asking of you?

OpenStudy (anonymous):

im not sure how to get it to cos (t) though

OpenStudy (anonymous):

using the equation editor

OpenStudy (anonymous):

write the first expression in terms of the second \[ \sec^2(t)\sin^2(t) \to \cos(t) \]

ganeshie8 (ganeshie8):

Yes, \(\large \sec^2t \sin^2t \) \(\large \frac{\sin^2t}{\cos^2t} \)

OpenStudy (anonymous):

but how do i change that in cos (t)

ganeshie8 (ganeshie8):

next, u need to change numerator to cos term ok

ganeshie8 (ganeshie8):

\(\sin^2t = 1 - \cos^2t\) simply use that in numerator

ganeshie8 (ganeshie8):

Yes, \(\large \sec^2t \sin^2t \) \(\large \frac{\sin^2t}{\cos^2t} \) \(\large \frac{1-\cos^2t}{\cos^2t} \) Now that everything in \(\cos\) terms, we're done.

OpenStudy (anonymous):

so its ok if its left as cos^2 (t) ?

ganeshie8 (ganeshie8):

yes thats what the question is asking us to do i think from the way i see it

OpenStudy (anonymous):

oh ok thank you so much!

ganeshie8 (ganeshie8):

np :)